# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v1

@@ Abstract
Semi-supervised semantic segmentation (SS-SS) aims to mitigate the heavy annotation burden of dense pixel labeling by leveraging abundant unlabeled images alongside a small labeled set. While current teacher-student consistency regularization methods achieve strong results, they often overlook a critical challenge: the precise delineation of object boundaries. In this paper, we propose BoundMatch, a novel multi-task SS-SS framework that explicitly integrates semantic boundary detection into the consistency regularization pipeline. Our core mechanism, Boundary Consistency Regularized Multi-Task Learning (BCRM), enforces prediction agreement between teacher and student models on both segmentation masks and detailed semantic boundaries. To further enhance performance and sharpen contours, BoundMatch incorporates two lightweight fusion modules: Boundary-Semantic Fusion (BSF) injects learned boundary cues into the segmentation decoder, while Spatial Gradient Fusion (SGF) refines boundary predictions using mask gradients, leading to higher-quality boundary pseudo-labels. This framework is built upon SAMTH, a strong teacher-student baseline featuring a Harmonious Batch Normalization (HBN) update strategy for improved stability. Extensive experiments on diverse datasets including Cityscapes, BDD100K, SYNTHIA, ADE20K, and Pascal VOC show that BoundMatch achieves competitive performance against state-of-the-art methods while significantly improving boundary-specific evaluation metrics. We also demonstrate its effectiveness in realistic large-scale unlabeled data scenarios and on lightweight architectures designed for mobile deployment.
@@ 1 Introduction
Semantic segmentation, the task of assigning a class label to every pixel, provides the detailed scene understanding crucial for many modern vision systems. It is particularly vital for autonomous driving, where precise identification of road boundaries, pedestrians, and vehicles is paramount for safe navigation [1, 2]. Beyond driving, it empowers robotic agents to interact intelligently with complex environments [3, 4], aids in clinical diagnostics through medical image analysis [5, 6], and enables large-scale environmental monitoring via remote sensing [7]. However, the reliance of high-performance models on vast amounts of dense, pixel-level annotations presents a major bottleneck, as manual labeling is extremely time-consuming and costly. Semi-supervised semantic segmentation (SS-SS) offers a compelling solution by aiming to achieve strong performance while drastically reducing annotation requirements, primarily leveraging large quantities of readily available unlabeled data.
Among various SS-SS strategies, Consistency Regularization (CR) has proven highly effective and is widely adopted, often implemented within teacher–student frameworks [8, 9]. In this setup, a “teacher” model (e.g., an EMA of the “student”) generates pseudo-labels on weakly perturbed unlabeled images, which then supervise the student model trained on strongly perturbed versions of the same images. This encourages the model to learn robust representations invariant to nuisances. While CR methods have significantly boosted overall segmentation accuracy (e.g., mIoU), they tend to focus on improving predictions for interior regions. Consequently, a critical challenge persists: the accurate delineation of object boundaries, which often remain imprecise. This limitation hinders reliability of downstream tasks by adding uncertainty (e.g. a pedestrian merging into the background). Notably, state-of-the-art fully supervised methods increasingly address this by incorporating boundary detection as an auxiliary multi-task learning (MTL) objective, yielding demonstrably sharper results [10, 11, 12].
Leveraging boundary information within the semi-supervised setting, however, remains largely underexplored. The few existing attempts either use boundary cues derived from segmentation pseudo-labels mainly to re-weight the consistency loss near edges [13, 14], or, like BoundaryMatch [15], introduce an auxiliary binary boundary detection task using single-stage features, reporting limited benefits when integrated with a modern CR method. Therefore, a crucial opportunity remains: to simultaneously enforce CR on both segmentation masks and detailed semantic boundary predictions derived from rich hierarchical features, and to explicitly fuse this boundary information back into the segmentation pipeline to create a synergistic learning dynamic within SS-SS.
@@
To fill this gap, we propose BoundMatch, a novel teacher–student framework designed to effectively integrate multi-task semantic boundary detection with CR for SS-SS. At its core, BoundMatch introduces Boundary Consistency Regularized Multi-Task Learning (BCRM), a mechanism that uniquely enforces prediction agreement between the teacher and student on both the segmentation masks and the semantic boundary maps extracted from hierarchical backbone features. This ensures consistent learning for regions and contours alike from unlabeled data. To maximize the utility of boundary information, BoundMatch incorporates two efficient fusion modules: Boundary-Semantic Fusion (BSF) injects learned boundary cues into the segmentation decoder, promoting more accurate contour localization in the final masks, while Spatial Gradient Fusion (SGF) refines the boundary predictions themselves by incorporating spatial gradients from the segmentation output, leading to higher-quality boundary pseudo-labels for the consistency objective. Built upon SAMTH, a strong baseline featuring Harmonious Batch Normalization (HBN) for enhanced stability, BoundMatch’s primary novelty lies in this synergistic boundary-centric consistency and fusion approach. It is also designed to be flexible and potentially combinable with other CR methods.
We validate BoundMatch through extensive experiments on diverse benchmarks including Cityscapes, BDD100K, SYNTHIA, ADE20K, and Pascal VOC. Results demonstrate that our method achieves competitive or state-of-the-art performance while significantly improving boundary-specific metrics. Furthermore, we show its effectiveness in realistic large-scale unlabeled data scenarios and its applicability to lightweight architectures suitable for deployment. Our main contributions are:
  • We introduce BoundMatch, the first SS-SS framework to apply consistency regularization to semantic boundaries within a multi-task teacher-student setup, enhanced by novel fusion modules (BSF and SGF) for synergistic learning.
  • We present SAMTH, a strong and stable SS-SS baseline incorporating Harmonious Batch Normalization (HBN), which benefits various CR methods.
  • We provide comprehensive experimental validation demonstrating BoundMatch’s state-of-the-art potential, significant boundary quality improvements, and practical utility across multiple datasets and realistic settings.
@@ 2 Related Work
Recognizing that segmentation accuracy is often hindered by imprecise object boundaries, boundary-aware methods have emerged. GSCNN [10] was among the first to achieve significant improvements by incorporating multi-task learning (MTL) of boundaries and segmentation. DecoupleSegNet [23] introduced specialized decoders for boundary and interior regions. Inspired by dedicated boundary detection methods [24, 25, 26], RPCNet [27] proposed an MTL framework integrating both tasks. CSEL [28] employed an affinity learning strategy to further refine segmentation. More recent work has focused on conditioning the backbone features with boundary information, exemplified by STDC [29] and SBCB [11]. MobileSeed [12] extended boundary-aware MTL to lightweight Transformer backbones suitable for mobile deployment. However, the application of MTL with semantic boundary detection within the SS-SS context remains largely unexplored, motivating our work.
@@
Fully supervised semantic segmentation methods require large amounts of costly pixel-wise annotations. SS-SS aims to mitigate this burden by leveraging abundant unlabeled data alongside a smaller labeled set. Consistency regularization (CR), which encourages models to produce consistent predictions under different input perturbations, is a cornerstone technique in SS-SS [30, 31, 32, 33, 34, 35, 36]. A common implementation involves the mean-teacher framework [8], where an exponential moving average (EMA) of the student model’s weights generates stable pseudo-labels for unlabeled data. CutMix-MT [8] integrated CutMix augmentation to enhance consistency. Subsequent works like AugSeg [37] and iMAS [38] explored stronger augmentation strategies. Other directions include using multiple teachers (DualTeacher [39]), incorporating contrastive learning (ReCo [40], U2PL [41]), and unifying various strong-to-weak perturbation regularizations (UniMatch [9] and its variants [42, 43, 44]). Recently, applying SS-SS principles to Transformer-based architectures has gained traction [45, 46, 47]. Our work builds upon these CR foundations by introducing boundary detection as an auxiliary task, an aspect largely missing in prior SS-SS literature. Recognizing the need to validate SS-SS in diverse, practical scenarios, we also evaluate our approach on lightweight architectures suitable for mobile deployment.
Boundary-Aware Semi-Supervised Semantic Segmentation.
While most SS-SS methods concentrate on pseudo-label refinement and consistency losses, explicit integration of boundary information remains rare. Graph-BAS3Net [5] utilized an edge detection auxiliary task for semi-supervised medical image segmentation. CFCG [13], one of the first approaches for natural images, proposed increasing the consistency loss weight for boundary regions identified from pseudo-label masks as training progresses. BoundaryMatch [15] introduced an auxiliary binary boundary detection task, similar to STDC [29], within an SS-SS framework. More recently, CW-BASS [14] proposed a boundary-aware self-training approach.
Although our work also employs boundary consistency, it differs significantly from BoundaryMatch [15]. We utilize a proven semantic boundary detection MTL approach leveraging hierarchical features [11], rather than binary boundaries from single-stage features. Crucially, we introduce explicit feature fusion mechanisms (BSF and SGF) to ensure the boundary information effectively enhances segmentation. As demonstrated in our experiments, this comprehensive integration yields substantial benefits, whereas BoundaryMatch reported minimal effectiveness from their boundary component in the SS-SS setting. Our work uniquely combines semantic boundary consistency regularization derived from hierarchical features with explicit fusion mechanisms, creating a synergistic approach specifically designed for improving boundary delineation in SS-SS.
@@ 3 Approach
In this section, we present BoundMatch, our multi-task framework that incorporates boundary detection into semi-supervised semantic segmentation (SS-SS). We begin by outlining the notation and preliminaries for SS-SS, including the teacher-student paradigm, in Secs. 3.1 and 3.2. We then introduce SAMTH, our baseline method detailed in Sec. 3.3. Building upon this baseline, Sec. 3.4 describes our core Boundary Consistency Regularized Multi-Task Learning (BCRM) framework. Subsequently, Sec. 3.5 details the complementary boundary-aware modeling and refinement strategies—Boundary-Semantic Fusion (BSF) and Spatial Gradient Fusion (SGF). Finally, Sec. 3.6 synthesizes these components (SAMTH, BCRM, BSF, and SGF) into the complete BoundMatch framework.
@@ 3.1 Notation and Preliminaries
with $p^{L}=f_{\theta}(x)$ representing the semantic segmentation prediction for the network.
3.2 Teacher–Student Framework for Consistency Regularization
The teacher–student framework, as shown in Fig. 2, is a popular paradigm for SS-SS. It employs consistency regularization (CR) between the predictions of a teacher network and a student network for unlabeled samples. The core idea is to encourage the student network to produce consistent predictions for an unlabeled image under different perturbations. The teacher network, which is typically an exponential moving average (EMA) of the student network, provides more stable and reliable predictions, which are then used to guide the student’s learning.
@@ 3.2 Teacher–Student Framework for Consistency Regularization
Specifically, given a weakly augmented image $u^{w}$ and its strongly augmented counterpart $u^{s}$, we first obtain the teacher’s prediction $p^{T,w}$ and generate hard pseudo-labels $\hat{p}^{T}=\operatorname{argmax}(p^{T,w})$. These hard pseudo-labels represent the teacher’s most confident predictions for each pixel in the unlabeled image.
@@
The student’s parameters are also updated using labeled set via $\mathcal{L}^{L}$. The teacher’s parameters $\theta^{T}$ are updated using EMA of the student’s parameters:
@@ 3.3 Strong Data Augmentation Mean-Teacher with Hard Pseudo-Labels (SAMTH)
Our baseline, SAMTH, builds upon the CutMix-MT framework [8] but distinguishes itself by using hard pseudo-labels, similar to recent methods such as AugSeg [37] and UniMatch [9]. SAMTH leverages CutMix as the strong data augmentation (SDA) to improve the robustness of the student network and the quality of the pseudo-labels.
A key differentiator of SAMTH is its Harmonious Batch Normalization (HBN) update strategy, addressing potential instabilities from BN handling in teacher-student frameworks. Common alternatives risk issues: forwarding only partial data through the teacher (e.g., CutMix-MT [8], U2PL [41]) can lead to biased BN statistics unsuitable for the teacher’s full role, while using the student’s BN statistics via copy or EMA (e.g., AugSeg [37], iMAS [38]) couples the teacher’s normalization to the student’s potentially noisy state and risks incompatibility between these borrowed BN statistics and the teacher’s own EMA-updated weights.
HBN mitigates these concerns. We forward the complete batch ($x,u^{w},u^{s}$) through both networks: $p^{S,L},,p^{S,w},,p^{S,s}=f_{\theta^{S}}(x,u^{w},u^{s})$ and $p^{T,L},,p^{T,w},,p^{T,s}=f_{\theta^{T}}(x,u^{w},u^{s})$. Crucially, the teacher’s BN layers remain in train mode, updating statistics independently based on the full data stream passing through the teacher’s own parameters. Subsequently, only the teacher’s non-BN weights are updated via EMA. This ensures the teacher’s BN statistics are harmonious with its specific weights and the full input distribution, decoupling its normalization from student fluctuations. This promotes more stable training dynamics and improves performance, as validated empirically across multiple baselines in our ablations (Sec. 4.3.5).

### v2

@@ Abstract
Semi-supervised semantic segmentation (SS-SS) aims to mitigate the heavy annotation burden of dense pixel labeling by leveraging abundant unlabeled images alongside a small labeled set. While current consistency regularization methods achieve strong results, they often overlook a critical challenge: the precise delineation of object boundaries. In this paper, we propose BoundMatch, a novel multi-task SS-SS framework that explicitly integrates semantic boundary detection into a teacher-student consistency regularization pipeline. Our core mechanism, Boundary Consistency Regularized Multi-Task Learning (BCRM), enforces prediction agreement between teacher and student models on both segmentation masks and detailed semantic boundaries. To further enhance performance and sharpen boundaries, BoundMatch incorporates two lightweight fusion modules: Boundary-Semantic Fusion (BSF) injects learned boundary cues into the segmentation decoder, while Spatial Gradient Fusion (SGF) refines boundary predictions using mask gradients, leading to higher-quality boundary pseudo-labels. This framework is built upon SAMTH, a strong teacher-student baseline featuring a Harmonious Batch Normalization (HBN) update strategy for improved stability. Extensive experiments on diverse urban-driving scene datasets including Cityscapes, BDD100K, and SYNTHIA show that BoundMatch achieves competitive performance against current state-of-the-art methods. Our approach achieves state-of-the-art results on the new benchmark with DINOv2 foundation model. We further validate our approach’s generalizability on Pascal VOC and ADE20K datasets. Ablation studies highlight BoundMatch’s ability to improve boundary-specific evaluation metrics, its effectiveness in realistic large-scale unlabeled data scenarios, and applicability to lightweight architectures for mobile deployment.
@@ 1 Introduction
Semantic segmentation, the task of assigning a class label to every pixel, provides the detailed scene understanding crucial for many modern vision systems. It is particularly vital for autonomous driving, where precise identification of road boundaries, pedestrians, and vehicles is paramount for safe navigation [1, 2]. Beyond driving, it empowers robotic agents to interact intelligently with complex environments [3, 4], aids in clinical diagnostics through medical image analysis [5, 6], and enables large-scale environmental monitoring via remote sensing [7]. However, the reliance of high-performance models on vast amounts of dense, pixel-level annotations presents a major bottleneck, as manual labeling is extremely time-consuming and costly. To address this issue, semi-supervised semantic segmentation (SS-SS) offers a compelling solution by aiming to achieve strong performance while drastically reducing annotation requirements, primarily leveraging large quantities of readily available unlabeled data.
Among various SS-SS strategies, Consistency Regularization (CR) has proven highly effective and is widely adopted, often implemented within teacher-student frameworks [8, 9]. In this setup, a “teacher” model (e.g. an exponential moving average (EMA) of the “student”) generates pseudo-labels on weakly perturbed unlabeled images, which then supervise the student model trained on strongly perturbed versions of the same images. This encourages the model to produce consistent predictions despite input perturbations. While CR methods have significantly boosted overall segmentation accuracy (e.g. mIoU), their improvements are primarily observed in interior regions, with boundary areas remaining challenging. This limitation hinders the reliability of downstream tasks by adding uncertainty (e.g. a pedestrian merging into the background). Notably, state-of-the-art fully supervised methods explicitly address this issue by incorporating boundary detection as an auxiliary multi-task learning (MTL) objective, yielding demonstrably sharper results [10, 11, 12].
Despite the proven benefits of boundary information in fully supervised settings, its integration into semi-supervised learning remains largely underexplored. Existing attempts fall into two categories: (1) methods that derive boundary cues from segmentation pseudo-labels to re-weight consistency losses near edges [13, 14], and (2) approaches like BoundaryMatch [15] that add binary boundary detection as an auxiliary task. The former relies on potentially noisy pseudo-label boundaries, while the latter treats boundary detection as a disconnected auxiliary task without explicit mechanisms to integrate boundary information back into segmentation, resulting in minimal performance gains. This reveals a crucial gap: the opportunity to enforce CR on semantic boundaries (rather than binary edges) derived from hierarchical features, while creating explicit fusion pathways that allow boundary and segmentation tasks to mutually enhance each other throughout the learning process.
@@
To fill this gap, we propose BoundMatch, a novel teacher-student framework that effectively integrates semantic boundary detection with CR for SS-SS. At its core, BoundMatch introduces Boundary Consistency Regularized Multi-Task Learning (BCRM), which enforces prediction agreement between teacher and student on both segmentation masks and semantic boundaries derived from hierarchical features. To create a synergistic relationship between tasks, we incorporate two fusion modules: Boundary-Semantic Fusion (BSF) injects boundary cues into the segmentation decoder, while Spatial Gradient Fusion (SGF) refines boundary predictions using spatial gradients, yielding higher-quality pseudo-labels. Additionally, we develop SAMTH, a strong teacher-student baseline that addresses training instabilities through Harmonious Batch Normalization (HBN), which maintains consistent batch normalization (BN) statistics between teacher and student networks. BoundMatch is designed as a flexible framework that can enhance various CR methods, as we demonstrate by integrating it with multiple baselines including SAMTH, UniMatch, and PrevMatch.
Extensive experiments demonstrate that BoundMatch achieves competitive or state-of-the-art performance across diverse benchmarks for urban-driving scenes, with particularly notable improvements in boundary quality metrics and strong results when using DINOv2 foundation models. BoundMatch could also generalize to other datasets such as ADE20K and Pascal VOC, and it maintains effectiveness in realistic large-scale unlabeled data scenarios.
Our main contributions are:
  • We provide a new perspective for SS-SS: that semantic boundaries can serve as an effective auxiliary task for consistency regularization. Through BCRM, we demonstrate that enforcing consistency on semantic boundaries (not just binary edges) derived from hierarchical features opens new possibilities for improving semi-supervised segmentation.
  • We propose BoundMatch, a comprehensive framework that realizes this perspective through careful architectural design. Beyond BCRM, BoundMatch incorporates two crucial fusion modules, BSF and SGF, that create bidirectional information flow between boundary and segmentation tasks. This synergistic design enables significant performance gains by allowing both tasks to mutually improve each other’s pseudo-labels.
  • We identify and resolve training instabilities in teacher-student frameworks using EMA update. Our HBN strategy yields SAMTH, a simple yet strong baseline that achieves competitive results even without boundary modeling, demonstrating the importance of proper BN handling in SS-SS.
  • We emphasize the importance of boundary-specific evaluation in SS-SS by systematically reporting BIoU and BF1 metrics alongside traditional mIoU, revealing that boundary quality improvements not captured by standard metrics.
  • Through extensive experiments across five datasets, we validate BoundMatch’s effectiveness when applied to various baselines, achieving state-of-the-art results with DINOv2 backbones, and demonstrating successful integration with lightweight architectures for real-world applications.
@@ 2 Related Work
Recognizing that segmentation accuracy is often hindered by imprecise object boundaries, boundary-aware methods have emerged. GSCNN [10] was among the first to achieve significant improvements by incorporating multi-task learning (MTL) of boundaries and segmentation. DecoupleSegNet [23] introduced specialized decoders for boundary and interior regions. Inspired by dedicated boundary detection methods [24, 25, 26], RPCNet [27] proposed an MTL framework integrating both tasks. CSEL [28] employed an affinity learning strategy to further refine segmentation. More recent work has focused on conditioning the backbone features with boundary information, exemplified by STDC [29] and SBCB [11]. Mobile-Seed [12] extended boundary-aware MTL to lightweight Transformer backbones suitable for mobile deployment. However, the application of MTL with semantic boundary detection within the SS-SS context remains largely unexplored, motivating our work.
@@
Fully supervised semantic segmentation methods require large amounts of costly pixel-wise annotations. SS-SS aims to mitigate this burden by leveraging abundant unlabeled data alongside a smaller labeled set. Consistency regularization (CR), which encourages models to produce consistent predictions under different input perturbations, is a cornerstone technique in SS-SS [30, 31, 32, 33, 34, 35, 36]. A common implementation involves the mean-teacher framework [8], where an exponential moving average (EMA) of the student model’s weights generates stable pseudo-labels for unlabeled data. CutMix-MT [8] integrated CutMix augmentation to enhance consistency. Subsequent works like AugSeg [37] and iMAS [38] explored stronger augmentation strategies. Teacher-Student framework has been further extended with various techniques, such as using multiple teachers (e.g. DualTeacher [39]) and contrastive learning (e.g. ReCo [40], U2PL [41]).
There are other notable directions in CR. Some methods focus on co-training to provide cross supervision by using two or more independent networks with different initializations or architectures (e.g. CPS [42] and Diverse Co-Training [43]). On the other hand, approaches like UniMatch [9] and its variants [44, 45, 46] have proposed a unified framework that combines strong-to-weak perturbation regularization with a single model training strategy.
Our work builds upon these CR foundations by introducing boundary detection as an auxiliary task, an aspect largely missing in prior SS-SS literature. Recently, applying SS-SS principles to Transformer-based architectures has gained traction [47, 48, 49, 50]. Recognizing the need to validate SS-SS in diverse, practical scenarios, we also evaluate our approach on DPT [51] with DINOv2 pretrained ViT backbones [52] and lightweight architectures suitable for mobile deployment.
Boundary-Aware SS-SS.
While most SS-SS methods concentrate on pseudo-label refinement and consistency losses, explicit integration of boundary information remains rare. Some approaches derive boundary masks from segmentation pseudo-labels using simple operators: CFCG [13] applies Laplacian operators while CW-BASS [14] uses Sobel filters. These masks are then used to re-weight the cross-entropy loss and put importance on boundary regions. While these methods demonstrate that boundary awareness can improve SS-SS, they are fundamentally limited to the boundary information already present in the pseudo-labels.
More relevant to our work are methods that explicitly incorporate boundary detection as an auxiliary task, learning boundaries independently rather than deriving them from segmentation outputs. Graph-BAS3Net [5] pioneered this direction for semi-supervised medical image segmentation. Most notably, BoundaryMatch [15] recently integrated binary boundary detection into SS-SS by adding an auxiliary boundary head atop UniMatch, following a design similar to STDC [29]. However, BoundaryMatch reported only marginal improvements when integrated with modern SS-SS methods, raising important questions about how to effectively leverage boundary detection in semi-supervised settings.
Our BoundMatch addresses these challenges through three key design decisions that differentiate it from BoundaryMatch: (1) We employ semantic (multi-label) boundaries rather than binary edges, providing richer class-specific supervision that better captures the nuances of different object categories; (2) We leverage hierarchical features from multiple backbone stages [11] instead of single-stage features, enabling multi-scale boundary reasoning; (3) Most critically, we introduce explicit fusion mechanisms (BSF and SGF) that create bidirectional information flow between boundary and segmentation tasks, rather than treating them as independent objectives. These design choices enable BoundMatch to achieve significant performance gains that we comprehensively measure using both standard mIoU and boundary-specific metrics (BIoU and BF1). Our systematic evaluation reveals that boundary-aware SS-SS methods provide substantial improvements in boundary quality that traditional metrics fail to capture, demonstrating that the key to effective boundary-aware SS-SS lies in creating true synergy between the auxiliary and main tasks rather than simply adding boundary detection as a disconnected component.
@@ 3 Approach
In this section, we present BoundMatch, our multi-task framework that incorporates boundary detection into semi-supervised semantic segmentation (SS-SS). We begin by outlining the notation and preliminaries for SS-SS, including the teacher-student paradigm, in Secs. 3.1 and 3.2. We then introduce SAMTH, our baseline method detailed in Sec. 3.3. Building upon this baseline, Sec. 3.4 describes our core Boundary Consistency Regularized Multi-Task Learning (BCRM) framework. Subsequently, Sec. 3.5 details the complementary boundary-aware modeling and refinement strategies: Boundary-Semantic Fusion (BSF) and Spatial Gradient Fusion (SGF). Finally, Sec. 3.6 synthesizes these components (SAMTH, BCRM, BSF, and SGF) into the complete BoundMatch framework.
@@ 3.1 Notation and Preliminaries
with $B^{L}$ representing mini-batch size and $p^{L}=f_{\theta}(x)$ representing the semantic segmentation prediction for the network.
3.2 Teacher-Student Framework for Consistency Regularization
The teacher-student framework, as shown in Fig. 2, is a popular paradigm for SS-SS. It employs consistency regularization (CR) between the predictions of a teacher network and a student network for unlabeled samples. The core idea is to encourage the student network to produce consistent predictions for an unlabeled image under different perturbations. The teacher network, which is typically an exponential moving average (EMA) of the student network, provides more stable and reliable predictions, which are then used to guide the student’s learning.
@@ 3.2 Teacher-Student Framework for Consistency Regularization
Specifically, given a weakly augmented image $u^{w}$ and its strongly augmented counterpart $u^{s}$, we first obtain the teacher’s prediction $p^{T,w}$ and generate hard pseudo-labels $\hat{p}^{T}=\operatorname{argmax}(p^{T,w})$. These pseudo-labels represent the teacher’s most confident predictions for each pixel in the unlabeled image.
@@
The student’s parameters are jointly updated using labeled set via $\mathcal{L}^{L}$. The teacher’s parameters $\theta^{T}$ are updated using EMA of the student’s parameters:
@@ 3.3 Strong Data Augmentation Mean-Teacher with Hard Pseudo-Labels (SAMTH)
Our baseline, SAMTH, builds upon the CutMix-MT framework [8] but distinguishes itself by using thresholded pseudo-labels (hard pseudo-labels) rather than Softmax probabilities (soft pseudo-labels), similar to recent methods such as AugSeg [37].
A key differentiator of SAMTH is its Harmonious Batch Normalization (HBN) update strategy, addressing potential instabilities in teacher-student training using Batch Normalization (BN) with EMA. Common alternatives risk issues:
  • Forwarding only partial data through the teacher (e.g. CutMix-MT [8], U2PL [41]) can lead to biased BN statistics unsuitable for the teacher’s full role.
  • Copying or taking the EMA of the student’s BN statistics (e.g. AugSeg [37], iMAS [38]) couples the teacher’s normalization to the student’s potentially noisy state and risks incompatibility between these borrowed BN statistics and the teacher’s own EMA-updated weights.
HBN aims to mitigate these concerns. We forward the complete batch ($x,u^{w},u^{s}$) through both networks: $p^{S,L},p^{S,w},p^{S,s}=f_{\theta^{S}}(x,u^{w},u^{s})$ and $p^{T,L},p^{T,w},p^{T,s}=f_{\theta^{T}}(x,u^{w},u^{s})$. Crucially, the teacher’s BN layers remain in train mode, updating statistics independently based on the full data stream passing through the teacher’s own parameters. For every iteration, the teacher updates its BN statistics, while the rest of the model parameters (i.e. layer weights) are EMA-updated from the student’s parameters. This ensures the teacher’s BN statistics are harmonious with its specific weights and the full input distribution, decoupling its normalization from student fluctuations. This promotes more stable training dynamics and improves performance, as validated empirically across multiple baselines in our ablations (Sec. 4.3.6). For more complete details and pseudocode, please refer to Appendix A.